OpenStudy (anonymous):

What rate of interest compounded quarterly will yield an effective interest rate of 5 %?

OpenStudy (kropot72):

If the effective yearly interest rate is 5%, the amount at the end of one year is the principal multiplied by 1.05. Let the yearly interest rate compounded quarterly be r. Then we can write the equation: \[\large (1+\frac{r}{4})^{4}=1.05\ ..........(1)\] Taking the fourth root of both sides of (1) we get: \[\large 1+\frac{r}{4}=\sqrt[4]{1.05}\ ........(2)\] which can be rearranged to give: \[\large r=4(\sqrt[4]{1.05}-1)=you\ can\ calculate\]

OpenStudy (anonymous):

Thank you, I got 12.40%

OpenStudy (anonymous):

Or 12.395% because it wants it rounded to 3 decimal places

OpenStudy (kropot72):

Unfortunately your result is way too large. Please post the steps in your calculation.

OpenStudy (anonymous):

I did it again and got .049. Not sure if that is any better though

OpenStudy (kropot72):

That is correct. To convert to a percentage multiply by 100: \[\large Percentage\ rate=0.049\times100=(you\ can\ calculate)\ percent\]